Pasha is thinking of a number such that when twice the number is added to three times one more than the number seh gets the same result as when she multiplies four times one less than the number. What number is Pasha thinking about?
Found 2 solutions by ankor@dixie-net.com, Ashley.hall:
Answer by ankor@dixie-net.com(22740)

Pasha is thinking of a number such that when twice the number is added to three
times one more than the number he gets the same result as when he multiplies four times one less than the number.
What number is Pasha thinking about?
:
Let x = "a number"
:
2x + 3(x+1) = 4(x-1)
2x + 3x + 3 = 4x - 4
5x - 4x = -4 - 3
x = -7 is the number
;
:
See if that works
2(-7) + 3(-7+1) = 4(-7-1)
-14 + 3(-6) = 4(-8)
-14 - 18 = -32

Answer by Ashley.hall(1)

2x + 3(x + 1) = 4(x - 1)
2x + 3x + 3 = 4x -4
5x +3 = 4x -4
-5x -5x
+3 = -1x -4
+4 +4
7 = -1x
___ ____
-1 -1
-7 = x